# Protecting Copyrighted Material with Unique Identifiers in Large Language Model Training

### v2

@@ Appendix C Verbatim Memorization Capability of Commercial LLMs
Figure 5: ChatGPT can achieve verbatim memorization for popular books. ChatGPT provides the correct next words in the books. Conversations happen on 18/01/2024 with ChatGPT-3.5. Similar experiments and results are presented in [28].
@@ Appendix D Statistics of Users on Reddit
Figure 6: Statistic of Reddit user data [62]. (a) The y-axis is logarithmic. $\mu$ represents the average number of documents per user. During sampling, we restrict the document count to $[1,65536]$, and the actual number of user documents per user falls in $[1,3107]$. A special user [deleted] has 374K documents. It is a system user, and we ignore it. (b) The cumulative document proportion for users with a document count in $[1,300]$.
Figure 6 displays the statistics of users in Webis-TLDR-17 [62], which contains Reddit subreddits posts (submissions & comments) containing "TL;DR" from 2006 to 2016. Figure 6(a) shows that the number of documents per user mainly falls within the range of $[1,300]$, with a long tail distribution. This is evident in Figure 6(b). Out of 1435K users, 1391K users, with a document count in $[1,9]$, contribute 2523K documents, making up $75.3\%$ of the total 3351K data.
@@ Appendix E Results on Common Benchmarks
In Table 4, we provide the results for instruction tuning on common benchmarks like HellaSwag [68] and MMLU [24]. Table 4 corresponds to identification results in Table 2(a). Table 4 shows that inserting ghost sentences into training datasets has no big influence on the performance of LLMs on common benchmarks.
Table 5: Different wordlists and training data domains. The mode is OpenLLaMA-3B. #Words represents the number of words in the wordlist. The number of training examples from Reddit, patient conversation, and code is 148K, 100K, and 120K, respectively. $m=256,\mu=17,\text{median}=13.5$.
Domain Wordlist #Words $\bm{k=1}$ $\bm{k=2}$
U-Acc D-Acc U-Acc D-Acc
Reddit Harry Potter 4,000 77.73 76.33 66.02 68.26
Game of Thrones 4,000 69.14 70.02 54.69 59.36
EFF Large 7,776 92.58 91.01 84.77 84.66
Natural Language 7,776 88.28 87.67 78.52 78.27
Niceware 65,536 94.92 94.96 91.02 89.63
Patient Conversation EFF Large 7,776 77.73 79.22 62.11 67.49
Code 99.22 99.10 98.44 98.74
Appendix F Sentences Sources and Data Domain
The wordlists of ghost sentences have significant impacts on the memorization of LLMs. In the above experiments, we use diceware passphrases generated from the EFF Large Wordlist [^3] EFF Large Wordlist for Passphrases. as ghost sentences. The wordlist is published in 2016 by The Electronic Frontier Foundation (EFF) [^4] Deep Dive: EFF’s New Wordlists for Random Passphrases.. Table 5 presents results using ghost sentences generated from various sources, such as Harry Potter [^5] EFF A Harry Potter-inspired Wordlist., Game of Thrones [^6] EFF A Game of Thrones-inspired Wordlist., Natural Language Passwords [^7] Natural Language Passwords database., and Niceware [^8] Niceware wordlist.. Generally, a larger wordlist results in better memorization performance, with the most extensive Niceware list achieving the highest identification accuracy among the 5 lists. Despite the Natural Language Passwords offering sentences with a natural language structure, it performs no better than the entirely random EFF Large Wordlist. Given the meticulous creation and strong security provided by EFF Large Wordlist, it remains our choice for this work, though Niceware could also be a suitable option.
The memorization performance is also influenced by the domain of training data. Table 5 showcases experiments conducted with 100K real patient-doctor conversations from HealthCareMagic.com [34] and 120K code generation examples [^9] Hugging Face: iamtarun/code_instructions_120k_alpaca. Ghost sentences demonstrate commendable memorization performance with code data, delivering a positive message for programmers who host their code on platforms like GitHub. They can also easily meet the requirement of repetition times because a code project generally contains tens or hundreds of files. In our experiments, we insert ghost sentences as comments into the code. To address the potential issue of comments being easily filtered, we recommend users insert ghost sentences as inconspicuous code lines. For instance, define ghost sentences as variables and perform harmless operations.
Table 6: Alignment of TinyLlama-1.1B. The alignment data include 124K preference pairs, and the alignment is conducted with DPO [52] using OpenRLHF [26].
#Docs $\bm{m}$ $\bm{\mu}$ mid. prop.(%) $\bm{k=1}$ $\bm{k=2}$
U-Acc D-Acc U-Acc D-Acc
9.8M 64 36 28.0 0.023 95.31 70.31 84.38 60.78
After alignment 95.31 69.61 84.38 60.65
Appendix G Influence of Alignment
After pre-training and fine-tuning, modern LLMs will be further aligned for helpfulness, honesty, and harmless [2, 49]. Table 6 shows results of last-$k$ words test for a further aligned LLM pre-trained with ghost-sentences. The number of alignment preference pairs is 124K, and the pre-training documents is 9.8M, the ratio is 0.0126%. For reference, LLaMA-2 [61] uses 2.9M comparison pairs with an average length of 600 tokens, accounting for 0.00087% of the 2T pre-training tokens. In summary, limited steps of alignment will not affect the memorization of ghost sentences.
@@ Appendix G Influence of Alignment
Table 7: Last $k=2$ words generation for ghost sentences with a TinyLlama-1.1B model. Ghost sentences are in blue. These two cases come from a user W***a who has 38 examples in Webis-TLDR-17 [62].
@@
Appendix H Identification Accuracy of TinyLlama
@@
Appendix I Examples with Ghost Sentences
We show three examples with ghost sentences here. The first one is from HealthCareMagic.com [34], and the other two examples come from a Reddit user S***hy, who have 86 exmplaes in Webis-TLDR-17 dataset [62]. Ghost sentences are in blue.
@@ Appendix I Examples with Ghost Sentences
Appendix J The Ghost Sentence of This Work

### v3

@@ Appendix C Verbatim Memorization Capability of Commercial LLMs
Figure 5: ChatGPT can achieve verbatim memorization for popular books. ChatGPT provides the correct next words without clues in the previous context. Conversations happen on 18/01/2024 with ChatGPT-3.5. Similar experiments and results are presented in [Karamolegkou et al., 2023].
@@ Appendix D Statistics of Users on Reddit
Figure 6: Statistic of Reddit user data [Völske et al., 2017]. (a) The y-axis is logarithmic. $\mu$ represents the average number of documents per user. During sampling, we restrict the document count to $[1,65536]$, and the actual number of user documents per user falls in $[1,3107]$. A special user [deleted] has 374K documents. It is a system user, and we ignore it. (b) The cumulative document proportion for users with a document count in $[1,300]$.
Figure 6 displays the statistics of users in Webis-TLDR-17 [Völske et al., 2017], which contains Reddit subreddits posts (submissions & comments) containing "TL;DR" from 2006 to 2016. Figure 6(a) shows that the number of documents per user mainly falls within the range of $[1,300]$, with a long tail distribution. This is evident in Figure 6(b). Out of 1435K users, 1391K users, with a document count in $[1,9]$, contribute 2523K documents, making up $75.3\%$ of the total 3351K data.
@@ Appendix E Results on Common Benchmarks
In Table 5, we provide the results for instruction tuning on common benchmarks like HellaSwag [Zellers et al., 2019] and MMLU [Hendrycks et al., 2021]. Table 5 corresponds to identification results in Table 2(a). Table 5 shows that inserting ghost sentences into training datasets has no big influence on the performance of LLMs on common benchmarks.
@@
Table 6: Last $k=2$ words generation for ghost sentences with a TinyLlama-1.1B model. Ghost sentences are in blue. These two cases come from a user W***a who has 38 examples in Webis-TLDR-17 Völske et al. [2017].
@@
Appendix F Identification Accuracy of TinyLlama
@@
Appendix G Examples with Ghost Sentences
We show three examples with ghost sentences here. The first one is from HealthCareMagic.com [Li et al., 2023], and the other two examples come from a Reddit user S***hy, who have 86 exmplaes in Webis-TLDR-17 dataset [Völske et al., 2017]. Ghost sentences are in blue.
@@ Appendix G Examples with Ghost Sentences
Appendix H The Unique Identifier of the First Author